OpenStudy (anonymous):

What is the exact value of the expressions the square root of 180. - the square root 125. + the square root of 5.? Simplify if possible. 2the square root of 2. 12the square root of 2. 2the square root of 5. 12the square root of 5.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Thanks by the way @ganeshie8 :)

ganeshie8 (ganeshie8):

simplify each term

ganeshie8 (ganeshie8):

\(\large \sqrt{180} = \sqrt{36\times 5} = ?\) \(\large \sqrt{125} = \sqrt{25\times 5} = ?\)